MATH 367, spg 98 PROJECT 1 – Due: Feb 9

You can work in groups of three but write the answers in your own words.

1. Make a summary with the 4 ways of counting learned in chapter 2.

2. Do problem 54 in chapter 2. Give a detailed explanation of your answers.

3. Do problem 55 in chapter 2. Give a detailed explanation of your answers.